# APVR: Hour-Level Long Video Understanding with Adaptive Pivot Visual Information Retrieval

### v1

@@ 3.3 Pivot Token Retrieval
Building upon the pivot frame selection mechanism, we propose a novel Pivot Token Retrieval framework that extends the concept of semantic importance from frame-level to token-level granularity. In this module, we adaptively retrieve the pivot visual token representations within the multimodal language model to maintain computational efficiency while preserving critical semantic information.
First, the retrieved pivot frames $\mathcal{F}_{pivot}$ are processed by a Vision Encoder $\mathcal{G}(\cdot)$ and a Projector $\mathcal{P}(\cdot)$ to generate dense visual token representations: $\mathcal{T}_{visual}=\mathcal{G}(\mathcal{P}(\mathcal{F}_{pivot}))=\{\mathcal{T}_{t}\mid\mathcal{T}_{t}\in R^{c\times d_{T}}\}_{t=1}^{K}$, where $d_{T}$ demotes the number of tokens per frame. Our objective is to retrieve the pivot token $\mathcal{T}_{pivot}$. To this end, we introduce two components: query-aware multi-layer attention scoring and adaptive token selection.
@@
The core innovation lies in leveraging the expanded semantic information from the LLM-augmented query to guide token selection across multiple transformer layers. This process captures both local and global semantic relationships between the visual content and the textual query. The attention scores for layer $l$ are computed as:
$A_{cross}^{(l)}=softmax(\frac{q_{text}^{(l)}\cdot(k_{visual}^{(l)})^{T}}{\sqrt{d}})$ (15)
, where $A_{cross}^{(l)}\in R^{B\times H\times d_{q}\times d_{v}}$, $q_{text}$ denotes the query states derived from the expanded semantic information $\{Q,D,M\}$ and $k_{visal}$ denotes the key states derived from the visual tokens $\{\mathcal{T}_{t}\}_{t=1}^{K}$. Then the token-level attention scores of each layer are aggregated at the query dimension:
$a^{(l)}=\sum_{d_{q}}A_{cross}^{(l)}$ (16)
This formulation captures how much attention each key token receives across all query positions, providing a comprehensive measure of its importance for the current layer’s computations.
@@
Dynamic Chunk-wise Selection: The long video understanding reaches the memory wall when prefilling the visual token. Thus, we perform token selection for the KV cache to reduce redundancy. Given a key or value cache to select, we first divide it into $W$ chunks at the token dimension. Then, the base selection ration of chunk $w$ is computed as:
$\eta_{w}^{(l)}=\frac{\sum_{w}a^{(l)}}{max(\{\sum_{w}a^{(l)}\mid w=1,...,W\})}$ (17)
The dynamic ratio each chunk is calculated by:
$\rho^{(l)}_{w}=min(1.0,\sqrt{\frac{|\{j:a_{j}^{(l)}>0.01\cdot max(a^{(l)})\}|}{L_{w}}})$ (18)
@@
$\gamma^{(l)}_{w}=\eta^{(l)}_{w}\cdot\rho^{(l)}_{w}$ (19)
Head-wise Soft Voting Based on the ratio, we can select pivot token in each chunk by:
$\mathcal{Z}_{w}^{(l)}=topk(a_{w}^{(l)}),k=\gamma_{w}^{(l)}\cdot L_{w}$ (20)
This results may appear significant disparities in different attention heads. To figure this, we introduce a head-wise voting for final token selection:
$\mathcal{Z}_{w}^{(l)}=topk(\sum_{j=1}^{h}softmax(a_{w,j}^{(l)})),k=\gamma_{w}^{(l)}\cdot L_{w}$ (21)
, where the softmax function calculate the per-head scores and give a soft voting via normalization and sum. Finally, we update the KV cache as selected pivot token in each layer:
$\mathcal{T}^{(l)}_{pivot}=Concat(\{\mathcal{Z}_{w}^{(l)}\}_{w=1}^{W})$ (22)
Finally, the selected pivot tokens will be fed into a MLLM to generate the final response.
@@ 4.1 Experimental setup
We evaluate the performance of APVR in two popular benchmarks that contains hour-level videos: LongVideoBench(Wu et al. 2024). A benchmark for long-context video understanding, consisting of 3,763 videos with 6,678 annotated multiple-choise questions across 17 categories of referred reasoning questions. It evaluates models across temporal reasoning, character tracking, and narrative understanding, making it particularly well-suited for evaluating APVR’s dual-granularity retrieval mechanisms in processing hour-long video content. VideoMME(Fu et al. 2024). A multi-modal evaluation benchmark focused on fine-grained video understanding tasks requiring precise visual-textual alignment and detailed semantic comprehension. It comprises 900 videos totaling 256 hours, with 2,700 manually labeled complex multiple-choice question-answer pairs across 30 subfields. It’s emphasis on both coarse-grained temporal understanding and fine-grained semantic detail extraction aligns precisely with APVR’s hierarchical information retrieval strategy. To highlight the importance of the pivot information retrieval for visual understanding, we do not use video subtitles for the two benchmarks.
@@
We integrate APVR with two baseline MLLMs, namely, Qwen2-VL-7B (Wang et al. 2024b) and Qwen2.5-VL-7B (Bai et al. 2025b). The MLLMs process a prompt involving the question, retrieved pivot video frames, and the multi-choice answer options to generate responses. We first densely extracted the raw frames from video at 2 fps. Then, the sampled frames are iteratively scored using two basic visual models: CLIP (Radford et al. 2021) with ViT-B-16 (Dosovitskiy et al. 2020) backbone, and Grounding-DINO (Liu et al. 2024b) with Swin-T (Liu et al. 2021) backbone, based on the expanded semantic information generated by LLM. The prompt for query expansion is in the supplementary materials. All evaluations are conducted on 8 NVIDIA A800 GPUs with 80GB memory, using LMMs-Eval (Zhang et al. 2024a) framework. Our APVR enables processing of up to 1024 frames, compared to the baseline 7B model’s capacity of 256 frames under optimal MLLM configurations.
@@ 4.2 Comparison to the State-of-the-Art
We first compare the accuracy of video question answering between APVR with recent training-based or training-free methods. Results are summarized in Table 1. Upon baseline of Qwen2-VL, APVR improves the overall accuracy by 9.5% and 3.0% in LongVideoBench and VideoMME, respectively. Upon a stronger baseline Qwen2.5-VL, our APVR also improves the overall accuracy by 8.3% in LongVideoBench and 4.6% in VideoMME, respectively. Notably, APVR based on Qwen2.5-Vl reaches 64.9 accuracy in LongVideoBench, which is higher than the powerful proprietary models such as GPT4-V and Gemini-1.5-Pro. The overall performance of APVR also surpasses other training-based and training-free methods with larger parameter scales such as 8B and 14B. Another notable result is that our APVR reaches highest accuracy on long and medium video samples of VideoMME, which demonstrates that our dual-granularity retrieval significantly assist capturing the most important information from extra-long videos and then improve the performance.
@@
In Figure 2, we illustrate representative video understanding results of APVR and the baseline Qwen2.5-VL, over four samples with hour-level video. First, it can be seen that the expanded semantic information significantly supplement various reference needed by subsequent retrieval. Next, it can be observed that the retrieved pivot frames are closely related to the query and the expanded semantic information. However, in many videos, there are many frames which are similar and all related to the text. Based on the frame score distribution, if we select only a few frames, the most pivot frames may be lost. Otherwise, if we select too many frames such as 1024 frames, we will face the memory wall. Based on APVR, one can select 1024 frames and then compress the visual token via our pivot token retrieval. As a result, MLLMs can get a comprehensive view of adequate and pivot content, and thus obtain the correct answer.
@@ 4.3 Ablation Studies
To prove the effectiveness of the two proposed components in APVR, we conduct an ablation study for PFR and PTR in LongVideoBench. The results are illustrated in Figure 3. It can be seen that, based on two different baseline MLLMs, the accuracies of APVR decline without both PFR and PTR. These results significantly demonstrate the rationality and validity of the designed dual-granularity pivot information retrieval strategy.
@@
First, we study the impact of retained frame number $k$ and the weight $\lambda$ of the two visual models for frame scoring. Results are summarized in Table 2. The ablation studies are conducted on LongVideoBench. It can be seen that the lager the frame number, the better the performance. This trend is similar in accuracy of all videos or long videos. For $\lambda$, we find that the best performance is achieved when we set a trade-off between the CLIP and Grounding-DINO model.
Second, we ablation the best setting of the core hyper-parameters, i.e., the iteration time $p$ and the initial sample stride $\nabla$ in the first iteration. It can be observed that the best performance is achieved when $p=3$ and $\nabla=4$. The accuracies of samples with long and medium videos decrease when $p$ and $\nabla$ are changed. This is because the adaptive frame scoring needs to matain a balance between sampling and distribution update.
@@ 5 Conclusion
We present APVR, a training-free framework that addresses the fundamental memory wall limitation in hour-level video understanding through adaptive hierarchical information retrieval. The framework’s dual-granularity approach, combining Pivot Frame Retrieval with Pivot Token Retrieval, enables processing of extensive video sequences while maintaining computational tractability and semantic accuracy. Experimental validation demonstrates substantial performance improvements across standard benchmarks, with APVR-enhanced models achieving state-of-the-art results among training-free approaches. The framework’s plug-and-play architecture ensures immediate deployment compatibility with existing MLLM systems without requiring expensive retraining procedures. This paradigm shift toward retrieval-based optimization provides a sustainable alternative to parameter scaling or extensive training.

### v2

@@ 3.3 Pivot Token Retrieval
Pivot Token Retrieval extends the concept of semantic importance from frame-level to token-level granularity. In this module, we adaptively retrieve the pivot visual token within MLLM to maintain computational efficiency while preserving critical semantic information.
First, the retrieved pivot frames $\mathcal{F}_{pivot}$ are processed by a Vision Encoder $\mathcal{G}(\cdot)$ and a Projector $\mathcal{P}(\cdot)$ to generate dense visual token representations: $\mathcal{T}_{vis}=\mathcal{G}(\mathcal{P}(\mathcal{F}_{pivot}))=\{\mathcal{T}_{t}\}_{t=1}^{K}$. Our objective is to retrieve the pivot token $\mathcal{T}_{pivot}$. To this end, we introduce two components: query-aware multi-layer attention scoring and adaptive token selection.
@@
We leverage the expanded semantic information from the LLM-augmented query to guide token selection across multiple transformer layers. First, the attention scores for layer $l$ are computed as:
$A_{cross}^{l}=\mathrm{softmax}(\frac{q_{text}^{l}\cdot(k_{vis}^{l})^{T}}{\sqrt{d}})$ (9)
, where $A_{cross}^{l}\in R^{h\times d\times d_{q}\times d_{v}}$, $q_{text}^{l}$ denotes the query states derived from $\{Q,Des,Sem\}$ and $k_{vis}^{l}$ denotes the key state derived from the visual tokens $\mathcal{T}_{vis}$. Then the token-level attention scores of each layer are aggregated at the query dimension by:
$a^{l}=\sum_{d_{q}}A_{cross}^{l}.$ (10)
This formulation captures how much attention each key token receives across all query positions, providing a comprehensive measure of its importance for the current layer’s computations. The superscript $l$ will be omitted for concision.
@@
Dynamic Chunk-wise Selection: The long video understanding reaches the memory wall when prefilling the visual token. Thus, we perform token selection for the KV cache to reduce redundancy. Given a key or value cache to select, we first divide it into $W$ chunks at the token dimension. Then, the base selection ratio of the chunk $w$ is computed as:
$\eta_{w}=\frac{\sum_{w}a}{\mathrm{max}(\{\sum_{i}a\mid i=1,...,W\})}$ (11)
The dynamic ratio of each chunk is calculated by:
$\rho_{w}=\mathrm{min}(1.0,\sqrt{\frac{|\{j:a_{j}>0.01\cdot\mathrm{max}(a)\}|}{L_{w}}})$ (12)
@@
$\gamma_{w}=\rho_{w}\cdot\eta_{w}$ (13)
Head-wise Soft Voting Based on the ratio $\gamma_{w}$, we can select the pivot token in each chunk by:
$\mathcal{Z}_{w}=\mathrm{TopK}(a_{w}\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (14)
This results may appear to be significant disparities in different attention heads. To figure out this, we introduce head-wise voting for final token selection:
$\mathcal{Z}_{w}=\mathrm{TopK}((\sum_{j=1}^{h}\mathrm{softmax}(a_{w,j}))\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (15)
, where the softmax function calculates the per-head scores and gives a soft voting via normalization and sum. Finally, we update the KV cache as the selected pivot token in each layer:
$\mathcal{T}_{pivot}=\mathrm{Concat}(\{\mathcal{Z}_{w}\}_{w=1}^{W})$ (16)
Finally, the selected pivot tokens will be fed into the MLLM to generate the final response.
@@ 4.1 Experimental setup
We evaluate the performance of APVR in two popular benchmarks that contains hour-level videos: LongVideoBench(Wu et al. 2024). A benchmark for long-context video understanding, consisting of 3,763 videos with 6,678 annotated multiple-choice questions across 17 categories of referred reasoning questions. VideoMME(Fu et al. 2024). A multi-modal evaluation benchmark focused on fine-grained video understanding tasks requiring precise visual-textual alignment and detailed semantic comprehension. It comprises 900 videos totaling 256 hours, with 2,700 manually labeled complex multiple-choice question-answer pairs across 30 subfields. To highlight the importance of the pivot information retrieval for visual understanding, we do not use video subtitles for the two benchmarks.
@@
We integrate APVR with two baseline MLLMs, namely, Qwen2-VL-7B (Wang et al. 2024b) and Qwen2.5-VL-7B (Bai et al. 2025b). The MLLMs process a prompt involving the question, retrieved pivot video frames, and the multi-choice answer options to generate responses. We first densely extracted the raw frames from video with $fps=2$. Then, the sampled frames are iteratively scored using two basic visual models: CLIP (Radford et al. 2021) with ViT-B-16 (Dosovitskiy et al. 2020) backbone, and Grounding-DINO (Liu et al. 2024b) with Swin-T (Liu et al. 2021) backbone, based on the expanded semantic information generated by LLM. All evaluations are conducted on 8 NVIDIA A800 GPUs with 80GB memory, using LMMs-Eval (Zhang et al. 2024a) framework. Under the above setup, our APVR enables processing of up to 1024 frames, compared to the baseline 7B model’s capacity of 256 frames under optimal MLLM configurations.
@@ 4.2 Comparison to the State-of-the-Art
We first compare the accuracy of video question answering between APVR with recent training-based or training-free methods. All results are summarized in Table 1. Upon the Qwen2-VL baseline, APVR improves the overall accuracy by 9.5% and 3.0% in LongVideoBench and VideoMME, respectively. Upon a stronger baseline Qwen2.5-VL, our APVR also improves the overall accuracy by 8.3% in LongVideoBench and 4.6% in VideoMME, respectively. Notably, APVR based on Qwen2.5-Vl reaches 64.9% accuracy in LongVideoBench, which is higher than the powerful proprietary models such as GPT4-V and Gemini-1.5-Pro. The overall performance of APVR also surpasses other training-based and training-free methods with larger parameter scales such as 8B and 14B. Another notable result is that our APVR reaches highest accuracy on long and medium video samples of VideoMME, confirming its unique capability to preserve critical spatio-temporal information in extended sequences.
@@
Fig. 3 presents representative qualitative results on hour-level video understanding, comparing our APVR framework with the baseline Qwen2.5-VL across four challenging samples. The results clearly demonstrate the unique advantages of APVR in both the semantic expansion and the subsequent retrieval process.
Firstly, the semantic information expansion module in APVR significantly enriches the original query, providing the downstream retrieval process with comprehensive object, attribute, and relational cues that are otherwise missing when using only the raw query. This richer semantic context enables the retrieval module to more accurately identify and prioritize frames that are truly relevant to the query intent.
Secondly, we observe that the pivot frames selected by APVR are highly correlated with both the query and the expanded semantic information, effectively filtering out irrelevant or redundant content. In long videos, there often exist temporally adjacent frames that are visually similar and all somewhat related to the query. If only a very limited number of frames are selected, it is easy to miss critical moments; conversely, retaining an excessive number of frames (e.g., 1024 frames) would cause out-of-memory (OOM) errors. APVR overcomes this dilemma by PFR and PTR and distills the most informative visual information within sufficient pivot frames. As a result, APVR empowers MLLM to efficiently comprehend long-form video content, providing accurate and detailed answers even in scenarios where baselines fail or generate hallucinations.
@@ 4.3 Ablation Studies
To systematically validate the effectiveness of the proposed dual-granularity retrieval strategy, we conduct ablation studies on the two core components of APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTR), using the LongVideoBench benchmark. The results, shown in Fig. 4, consistently demonstrate that removing either PFR or PTR leads to a notable decline in accuracy across both baseline MLLMs. Without PFR, the system loses its ability to precisely localize query-relevant frames, resulting in performance degradation due to the inclusion of irrelevant or redundant content. Without PTR, the model is less able to compress visual tokens and focus on the most informative visual details, leading to processing maximum to only 256 frames and reduced answer quality. These findings strongly support the necessity and rationality of our dual-granularity pivot information retrieval design.
@@
We further conduct a comprehensive analysis of key hyper-parameters that influence the performance of APVR. First, we examine the impact of the number of retrieved pivot frames $K$ and the weighting parameter $\lambda$ that balances the contributions of the two visual models in frame scoring. As summarized in Table 2, increasing the number of retained frames generally improves performance, especially on long videos, by reducing the risk of missing relevant segments. For the model weighting parameter $\lambda$, we observe that optimal results are achieved by striking a balance between CLIP-based semantic similarity and Grounding-DINO-based object detection, confirming the complementarity of these two perspectives.
Second, we analyze the effect of core retrieval hyper-parameters, specifically the number of retrieval iterations $p$ and the initial sampling stride $\nabla$. As shown in Table 3, the best performance is observed when $P=3$ and $\nabla=4$, indicating that a moderate number of adaptive iterations and a balanced initial sampling density allow the system to progressively refine frame selection without causing overfitting or excessive computational cost. Deviating from these settings leads to decreased accuracy, highlighting the importance of carefully tuning the retrieval process to maintain a balance between sampling and distribution update.
@@ 5 Conclusion
We propose APVR, a training-free framework that overcomes the memory wall in hour-level video understanding via adaptive, hierarchical visual information retrieval. By integrating Pivot Frame Retrieval and Pivot Token Retrieval, APVR enables efficient and accurate processing of long videos without sacrificing semantic detail. Extensive experiments on standard benchmarks demonstrate that APVR delivers state-of-the-art results among training-free and training-based methods. Its plug-and-play design allows seamless integration with existing MLLM systems, eliminating the need for costly retraining. Our results establish retrieval-based optimization as a practical and scalable solution for long-form video comprehension.